OpenStudy (anonymous):

Solve the following system of equations. Please show your work to receive full credit. x - y = 10 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

you can add the two equations together (one option) which is easy because the top is -y and the bottom is +y, so they cancel when you add them, and you get 2x+x=10+2, so 3x=12. solve for x, then plug back in to ONE of the equations (doesn't matter which), and solve for y

OpenStudy (anonymous):

Thanks :)